OpenStudy (anonymous):

What is the y-intercept of the equation 3y + 2x = 12? A.(0,4) B.(4,0) C.(4,4) D.(0,0)

OpenStudy (anonymous):

Look : If we want to have a standard form should be have y alone . So have this : y=mx+d So y is alone means nothing is before or after that ... So we have3y + 2x = 12= 3y=-2x+12 But 3 is extra .. So delete it from this way ...:(Divided on 3) y=-2x/2+4 So y-intercept is 4 .! Got it ?:) @dom30012345123

OpenStudy (anonymous):

ok so its b.(4,0)?

OpenStudy (anonymous):

@sione

OpenStudy (anonymous):

We have : In the b : 3*4+0=12 So it s right :) Yea !

OpenStudy (anonymous):

ok thank you

OpenStudy (texaschic101):

to find the y intercept, you could just sub in 0 for x 3y + 2x = 12 3y + 2(0) = 12 3y = 12 y = 4 no....y intercept is (0,4) not (4,0) because your solution set is (x,y) answer is A

OpenStudy (anonymous):

thank you!

OpenStudy (texaschic101):

no problem :)